# An Evolutionary Algorithm to Accelerate High-Dimensional Optimization

### v2

@@ 3.4. Statistical Analysis:
The ratios of run times $R_{t}$ for QUASAR vs. DE & L-SHADE are shown in Table 7 & Table 8, by dimension and sample size respectively. $R_{t}$ is calculated by dividing the average run time of the comparison method ($\text{Mean Sec}_{\text{comparison}}$) by the average run time of QUASAR ($\text{Mean Sec}_{\text{QUASAR}}$) across all trials for that specific dimension.
(11) $R_{t}=\frac{\text{Mean Run Time}_{\text{comparison},D}}{\text{Mean Run Time}_{\text{QUASAR},D}}$
The overall efficiency metrics are the overall run time ratios, $R_{t_{overall}}$, calculated as the arithmetic mean of each paired ratio $R_{t}$.
(12) $\text{R}_{\text{t,overall}}=\frac{1}{|D|}\sum_{D\in D}\left(\frac{1}{|R_{t,D}|}\sum_{i=1}^{|R_{t,D}|}\frac{t_{\text{comparison},i,D}}{t_{\text{QUASAR},i,D}}\right)$
Statistical Significance:
For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is used. This non-parametric test was chosen as the final error distributions are on different scales for each function and are assumed not to be normal, and to ensure consistency with the non-parametric ranking methodology.
The Wilcoxon test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
@@ 4. Results
The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $GMERF$, quantifying QUASAR’s improvements in final solution fitness. The computational efficiency is also measured by comparing full optimization run times.
Since lower errors and run times indicate better performance, a $\text{ratio}>1$, for both the $GMERF$ and run times ratios, indicates that QUASAR demonstrated superior performance.
Similarly, all plots shown are of raw fitness errors and run times; for all plots, lower values indicate higher performance.
@@ 4.1. Solution Quality
The ranks and geometric means of fitness error $GMERF$ of the final solution are calculated for both the dimension-variant and sample-variant trials.
The 95% confidence intervals show consistent performance improvements in almost every scenario tested.
4.1.1. Dimension-Variant Trials:
The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions ($10D$, $30D$, $50D$, and $100D$), using a fixed sample size $N=\text{1000}$ and maximum generations $g_{max}=100$.
The geometric mean error reduction factors ($GMERF$) were consistently high across all dimensions (Fig 3). The 95% confidence intervals validate its consistent improvements; the intervals do not fall below $1.0\times$ GMERFGMERF.
The largest improvements were seen against DE in $100D$, achieving a $GMERF$ of $\text{13.52}\times$ with a CI upper bound of $16.68\times$. The comparison against L-SHADE showed similar improvements; the full set of $GMERF$ values are shown in Table 5.
@@ 4.1.1. Dimension-Variant Trials:
The Friedman test was used on the entire set of unique function/dimension combinations to determine the overall performance hierarchy of the optimizers. The test confirmed a statistically significant improvement in performance ($p\ll 0.001$), shown in Table 3.
Table 3. Dimension-Variant Rank Sums
Method Rank Sum
QUASAR 150
L-SHADE 229
DE 305
4.1.2. Sample Size-Variant Trials:
@@
The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $GMERF$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE.
These performance increases scale super-linearly with population size (Fig. 4, Table 6). The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ (low-population) case, where L-SHADE achieved a $GMERF$ of $0.49\times$.
The Friedman test is run on the set of all function/population size combinations, shown in Table 4.
@@
Table 4. Sample-Variant Rank Sums
Method Rank Sum
QUASAR 217
L-SHADE 223
DE 430
Table 5. Dimension-Variant GMERF Results: QUASAR vs L-SHADE and DE
D vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
10 1.30 (1.20, 1.40) 1.06 (1.01, 1.12) $7\times 10^{-16}$ $7\times 10^{-3}$ $1.59\times 10^{3}$ $2.10\times 10^{3}$ $1.72\times 10^{3}$
30 4.14 (3.62, 4.74) 4.94 (4.24, 5.77) $5\times 10^{-91}$ $3\times 10^{-108}$ $5.10\times 10^{3}$ $2.11\times 10^{4}$ $2.52\times 10^{4}$
50 3.78 (3.32, 4.31) 8.44 (7.09, 10.04) $9\times 10^{-79}$ $5\times 10^{-134}$ $1.46\times 10^{4}$ $5.51\times 10^{4}$ $1.23\times 10^{5}$
100 2.67 (2.32, 3.06) 13.52 (10.95, 16.68) $2\times 10^{-36}$ $2\times 10^{-141}$ $7.917\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
Table 6. Sample Size-Variant GMERF Results: QUASAR vs L-SHADE & DE
S vs L-SHADE vs DE p-value p-value QUASAR L-SHADE DE
GMERF 95% CI GMERF 95% CI (vs L-SHADE) (vs DE) (GM Error) (GM Error) (GM Error)
100 0.49 (0.44, 0.56) 2.43 (2.04, 2.89) $7\times 10^{-86}$ $3\times 10^{-46}$ $5.94\times 10^{5}$ $2.94\times 10^{5}$ $1.44\times 10^{6}$
250 1.34 (1.19, 1.50) 7.86 (6.51, 9.49) $3.3\times 10^{-2}$ $3\times 10^{-141}$ $1.79\times 10^{5}$ $2.39\times 10^{5}$ $1.41\times 10^{6}$
500 1.93 (1.70, 2.20) 10.40 (8.54, 12.67) $2\times 10^{-8}$ $2\times 10^{-140}$ $1.15\times 10^{5}$ $2.21\times 10^{5}$ $1.19\times 10^{6}$
1000 2.66 (2.32, 3.06) 13.52 (10.95, 16.68) $5\times 10^{-34}$ $2\times 10^{-141}$ $7.92\times 10^{4}$ $2.11\times 10^{5}$ $1.07\times 10^{6}$
5000 4.47 (3.41, 5.86) 18.54 (12.49, 27.53) $3\times 10^{-37}$ $3\times 10^{-48}$ $4.20\times 10^{4}$ $1.88\times 10^{5}$ $7.78\times 10^{5}$
@@ 4.2. Computational Efficiency
The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{t_{overall}}$ (seconds) relative to DE and L-SHADE.
4.2.1. Dimension-Variant Trials:
In the dimension-variant trials, QUASAR demonstrated consistently faster run times compared to DE and L-SHADE.
The overall average run time improvement ratios $R_{t_{overall}}$ were $1.42\times$ and $6.29\times$ respectively. The average paired run time ratios $R_{t}$ are shown in Table 7, and optimization run times in seconds are in Table 7 and Fig. 5.
@@ 4.2.1. Dimension-Variant Trials:
4.2.2. Sample Size-Variant Trials:
In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig. 6, Table 8).
The $N=250$ trials, however, demonstrated a ratio $R_{t}$ vs DE of $0.84\times$, indicating that DE was approximately 16% faster.
@@ 4.2.2. Sample Size-Variant Trials:
Table 7. Dimension-Variant Run Times: QUASAR vs L-SHADE & DE
D $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
10 $4.21$ $1.26$ $5\text{\times}{10}^{-144}$ $7.3\text{\times}{10}^{-141}$ $16.4$ $69.1$ $20.7$
30 $3.17$ $1.17$ $5\text{\times}{10}^{-144}$ $1.1\text{\times}{10}^{-139}$ $25.0$ $79.4$ $29.2$
50 $2.66$ $1.14$ $5\text{\times}{10}^{-144}$ $1.6\text{\times}{10}^{-137}$ $33.5$ $89.0$ $38.1$
100 $2.03$ $1.09$ $5\text{\times}{10}^{-144}$ $3\text{\times}{10}^{-137}$ $55.4$ $112.4$ $60.5$
Table 8. Sample Size-Variant Run Times: QUASAR vs L-SHADE & DE
S $\mathbf{R_{t}}$ $\mathbf{R_{t}}$ p-value p-value QUASAR L-SHADE DE
$\mathbf{vs\ \textbf{L-SHADE}}$ $\mathbf{vs\ DE}$ (vs L-SHADE) (vs DE) (sec) (sec) (sec)
100 $1.37$ $1.05$ $\text{\times}{10}^{-142} 1×1$$\text{\times}{10}^{-84}$ $5.8$ $7.9$ $6.0$
250 $1.47$ $0.84$ $1\text{\times}{10}^{-142}$ $5\text{\times}{10}^{-58}$ $13.7$ $20.1$ $11.4$
500 $1.65$ $1.06$ $1\text{\times}{10}^{-142}$ $2\text{\times}{10}^{-123}$ $28.7$ $47.2$ $30.5$
1000 $2.00$ $1.08$ $1\text{\times}{10}^{-142}$ $9\text{\times}{10}^{-121}$ $57.3$ $114.4$ $62.0$
5000 $5.96$ $1.29$ $3\text{\times}{10}^{-49}$ $3\text{\times}{10}^{-48}$ $292.0$ $1741.7$ $376.1$
@@ 5. Ease of Use and Implementation
QUASAR is designed to be robust with minimal parameter dependence, making it effective for most numerical optimization problems.
@@ 5.1. Ease of Use
A key feature of QUASAR is its ability to operate effectively with its default entanglerateentangle\ rate, bypassing the hyperparameter tuning often required by other metaheuristics such as DE and L-SHADE.
The algorithm relies on its quasi-adaptive mechanisms to dynamically balance exploration and exploitation. For typical use cases, the only inputs needed are the objective function, search bounds, and the target number of generations ($g_{\max}$).
@@ 5.2. Implementation
QUASAR is implemented as part of a streamlined open-source Python package, designed for seamless integration into existing workflows. The syntax follows the structure of SciPy’s optimization module, making it nearly one-to-one interchangeable with its Differential Evolution.
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package containing QUASAR, hdim-opt, is available via the Python Package Index (PyPI) using ’pip install hdim_opt’, or at github.com/jgsoltes/hdim-opt which includes the experimental notebooks for full reproducibility.
@@ 6. Conclusion
This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit high-dimensional spaces.
The algorithm utilizes a highly stochastic architecture, including probabilistic mutation mechanisms, rank-based crossover rates, and asymptotically decaying covariance reinitializations.
Experimental trials across the CEC2017 benchmark suite confirm that QUASAR delivers exceptional performance compared to Differential Evolution (DE) and L-SHADE, consistently ranking best via the Friedman test.
[…]
The computational efficiency of the algorithm is a key benefit, with run times consistently lower than the comparison optimizers ($R_{t}>1$), scaling positively with population size and dimensionality.
@@ 7.1. Enhancements:
The current covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equall;. future work will investigate applying weights to the very best solutions. New methods for defining the probability of occurrence may be explored, such as sinusoidal or sigmoid activation behavior.
Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach, similar to the other QUASAR components, via soft-max or similar methods.
7.2. Verification:
QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems. The algorithms’s robustness would be further validated by rigorously evaluating its performance on more constrained, non-parametric problems.

### v3

@@ 3.4.3. Computational Efficiency:
The run times ratios $R_{\text{time}}$ for QUASAR vs. DE & L-SHADE are shown in Table 5, by dimension and population size. $R_{\text{time}}$ is calculated by dividing the arithmetic average run time $\bar{t}$ of the comparison method by the average run time of QUASAR across all trials for that specific dimension.
$R_{\text{time}}=\frac{\bar{t}_{\text{comp}}}{\bar{t}_{\text{QUASAR}}}=\frac{\sum_{i=1}^{T}t_{\text{comp},i}}{\sum_{i=1}^{T}t_{\text{QUASAR},i}}$
The overall efficiency metric, $\bar{R}_{\text{time}}$, is defined as the geometric mean of the $R_{\text{time}}$ ratios across all $k$ function/dimension or function/sample size combinations.
(11) $\bar{R}{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$
3.4.4. Statistical Significance:
For direct comparison within each dimension or sample size, the paired Wilcoxon signed-rank test is performed. This non-parametric test was chosen as the final error distributions are on different scales for each function, are assumed not to be normally distributed, and to ensure consistency with the non-parametric ranking methodology.
The signed-rank test was conducted on the final error values ($f$) and run times to generate the p-values for each function/dimension combination.
@@ 4. Results
The primary solution quality metrics are the Friedman aggregated rank sum and the geometric mean error reduction factor $R_{\text{error}}$, quantifying QUASAR’s improvements in final solution fitness. The computational efficiency is measured by comparing the ratio of full optimization run times $R_{\text{time}}$.
QUASAR demonstrated top performance in the rank sum test, shown in Table 2. Similarly, it demonstrated consistent improvements in geometric mean error reduction factor $R_{\text{error}}$ (Table 4) and run times $R_{\text{time}}$ (Table 5).
Table 2. Rank Sums by Sample Size and Dimension
N D Overall
QUASAR 217 150 367
L-SHADE 223 229 452
DE 430 305 735
Table 3. Overall Performance Comparison (Geometric Mean Ratios)
vs L-SHADE vs DE
$\mathbf{\bar{R}}$ 95% CI p-value $\mathbf{\bar{R}}$ 95% CI p-value
Quality ($\mathbf{\bar{R}_{\text{error}}}$) 2.07 (0.62, 6.96) $5.30\times 10^{-144}$ 3.85 (0.93, 16.03) $5.78\times 10^{-136}$
Run Time ($\mathbf{\bar{R}_{\text{time}}}$) 5.16 (4.57, 5.83) $3.36\times 10^{-141}$ 1.40 (1.36, 1.44) $4.52\times 10^{-23}$
Ratios ¿ 1 indicate superior QUASAR performance.
@@ 4.1. Solution Quality
The ranks and geometric mean fitness errors $R_{\text{error}}$ of the final solutions are calculated for both the dimension-variant and sample-variant trials. Solution quality comparisons are shown in Table 4.
4.1.1. Dimension-Variant:
The dimension-variant analysis assessed QUASAR’s convergence against DE and L-SHADE across four discrete dimensions $[10,30,50,100]$, using a fixed sample size $N=1000$ and maximum generations $g_{max}=100$.
The geometric mean error reduction factors ($R_{\text{error}}$) were consistently high across all dimensions (Fig. 3). The 95% confidence intervals validate its consistent improvements; the intervals do not fall below $R_{\text{error}}=1.0\times$. The largest improvements were seen against DE in $100D$, achieving a $R_{\text{error}}$ of $\text{13.52}\times$ with a CI upper bound of $16.68\times$. The comparison against L-SHADE showed the highest improvements in $30D$, with $R_{\text{error}}=4.14$, reaching CI upper bounds of 4.74.
@@ 4.1.1. Dimension-Variant:
4.1.2. Sample-Variant:
@@
The fitness improvement factors saw the highest gains with higher population sizes; in the $N=5000$ case, QUASAR reached a geometric mean error improvement factor $R_{\text{error}}$ of $\text{18.5}\times$ compared to DE, and $4.47\times$ compared to L-SHADE. These performance increases scale with population size (Fig. 4). The only scenario where QUASAR did not outperform L-SHADE was the $N=100$ case, where L-SHADE achieved $R_{\text{error}}=0.49\times$, highlighting the impact of bypassing the covariance reinitialization for $N<D$.
@@
Table 4. Solution Quality Comparison (Geometric Mean Error)
vs L-SHADE vs DE
$\mathbf{R_{error}}$ 95% CI p-value $\mathbf{R_{error}}$ 95% CI p-value
$\mathbf{D=10}$ 1.30 (1.20, 1.40) $7.0\times 10^{-16}$ 1.06 (1.01, 1.12) $7.0\times 10^{-3}$
$\mathbf{D=30}$ 4.14 (3.62, 4.74) $5.0\times 10^{-91}$ 4.94 (4.24, 5.77) $3.0\times 10^{-108}$
$\mathbf{D=50}$ 3.78 (3.32, 4.31) $9.0\times 10^{-79}$ 8.44 (7.09, 10.04) $5.0\times 10^{-134}$
$\mathbf{D=100}$ 2.67 (2.32, 3.06) $2.0\times 10^{-36}$ 13.52 (10.95, 16.68) $2.0\times 10^{-141}$
$\mathbf{N=100}$ 0.49 (0.44, 0.56) $7.0\times 10^{-86}$ 2.43 (2.04, 2.89) $3.0\times 10^{-46}$
$\mathbf{N=250}$ 1.34 (1.19, 1.50) $3.3\times 10^{-2}$ 7.86 (6.51, 9.49) $3.0\times 10^{-141}$
$\mathbf{N=500}$ 1.93 (1.70, 2.20) $2.0\times 10^{-8}$ 10.40 (8.54, 12.67) $2.0\times 10^{-140}$
$\mathbf{N=1000}$ 2.66 (2.32, 3.06) $5.0\times 10^{-34}$ 13.52 (10.95, 16.68) $2.0\times 10^{-141}$
$\mathbf{N=5000}$ 4.47 (3.41, 5.86) $3.0\times 10^{-37}$ 18.54 (12.49, 27.53) $3.0\times 10^{-48}$
@@ 4.2. Computational Efficiency
The computational efficiency of QUASAR was measured using the overall ratio of run time $R_{\text{time}}$ (seconds) relative to DE and L-SHADE. The full set of comparisons is shown in Table 5.
4.2.1. Dimension-Variant:
In the dimension-variant trials, QUASAR demonstrated consistently faster run times. The results, shown in Fig. 5, consistently show that QUASAR achieved faster optimization speeds across all dimensions tested, scaling linearly with dimensionality.
@@ 4.2.1. Dimension-Variant:
4.2.2. Sample-Variant:
In the sample-variant trials, run times were consistently faster than DE and L-SHADE, averaging $7.78\times$ and $1.39\times$ speed increases respectively across all sample-variant trials (Fig. 6).
The $N=250$ trials, however, demonstrated a ratio $R_{\text{time}}$ vs. DE of $0.84\times$, indicating that DE was approximately 16% faster. This likely represents the overhead of the covariance matrix calculation relative to the low total evaluation count, whereas the covariance calculations are bypassed in the lower $N=100$ case where QUASAR was faster than DE.
@@ 4.2.2. Sample-Variant:
Table 5. Efficiency Comparison (Geometric Mean Run Time)
vs L-SHADE vs DE GM Run Time (s)
$\mathbf{R_{\text{time}}}$ p-value $\mathbf{R_{\text{time}}}$ p-value QUASAR L-SHADE DE
$\mathbf{D=10}$ 4.21 $5\times 10^{-144}$ 1.26 $7.3\times 10^{-141}$ 16.4 69.1 20.7
$\mathbf{D=30}$ 3.17 $5\times 10^{-144}$ 1.17 $1.1\times 10^{-139}$ 25.0 79.4 29.2
$\mathbf{D=50}$ 2.66 $5\times 10^{-144}$ 1.14 $1.6\times 10^{-137}$ 33.5 89.0 38.1
$\mathbf{D=100}$ 2.03 $5\times 10^{-144}$ 1.09 $3\times 10^{-137}$ 55.4 112.4 60.5
$\mathbf{N=100}$ 1.37 $1\times 10^{-142}$ 1.05 $1\times 10^{-84}$ 5.8 7.9 6.0
$\mathbf{N=250}$ 1.47 $1\times 10^{-142}$ 0.84 $5\times 10^{-58}$ 13.7 20.1 11.4
$\mathbf{N=500}$ 1.65 $1\times 10^{-142}$ 1.06 $2\times 10^{-123}$ 28.7 47.2 30.5
$\mathbf{N=1000}$ 2.00 $1\times 10^{-142}$ 1.08 $9\times 10^{-121}$ 57.3 114.4 62.0
$\mathbf{N=5000}$ 5.96 $3\times 10^{-49}$ 1.29 $3\times 10^{-48}$ 292.0 1741.7 376.1
@@ 5.1. Ease of Use
QUASAR operates most effectively with its default parameters, bypassing the hyperparameter tuning often required by other metaheuristics. The algorithm relies on its quasi-adaptive mechanisms to dynamically balance exploration and exploitation.
Typical use cases require only the objective function, search bounds, and the target number of generations ($g_{\max}$). For additional control, the entangle rate can be tuned between $[0,1]$, though the default value of 0.33 was most robust across the CEC benchmarks.
While not used in this study, the optimizer naturally handles vectorized objective functions by evolving the full population at once. This can lead to orders of magnitude faster run times, depending on population size.
@@ 5.2. Implementation
QUASAR is implemented as part of a streamlined open-source Python package, designed for ease of use and adoption into existing workflows. The syntax follows the structure of SciPy’s optimize module, making it nearly one-to-one interchangeable with its Differential Evolution function.
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package containing QUASAR, hdim_opt, is available via PyPi or at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
@@ 6. Conclusion
This study introduced Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), a quantum-inspired evolutionary algorithm designed to efficiently explore and exploit complex high-dimensional spaces.
Combining probabilistic mutation, rank-based crossover, and asymptotically decaying covariance reinitializations, QUASAR’s highly stochastic architecture is exceptionally suited for discontinuous and non-parametric optimization. Its reliance on default parameters and streamlined implementation allows for seamless integration within existing workflows.
Experimental trials across the CEC2017 benchmark suite confirm that QUASAR delivers exceptional performance, achieving the highest rank via the Friedman test compared to Differential Evolution (DE) and L-SHADE, with respective overall geometric mean error improvements of $3.85\times$ and $2.07\times$.
QUASAR also proves computationally efficient, with overall geometric mean run time improvements of $1.40\times$ and $5.16\times$ over DE and L-SHADE. Analysis by dimension and population size shows linear improvement scaling across all metrics.
@@ 7.1. Enhancements
7.1.1. Mutations:
The mutation factors are sampled from fixed Gaussian distributions; dynamic scaling based on convergence may significantly help to maintain a healthy balance of exploration and exploitation. The strategies may benefit from different permutations of equation terms, or non-probablistic selction methods.
7.1.2. Crossover:
Crossover rate is currently defined by fitness rank, and may benefit from a probability-based approach similar to the other QUASAR components, via soft-max or similar methods. Many techniques from existing evolutionary frameworks may prove beneficial to the model.
7.1.3. Reinitialization:
The current reinitialization covariance matrix ($\mathbf{\Sigma}$) treats all top solutions equally; future work will investigate weighted solutions similar to those found in CMA-ES. New methods for defining the probability of occurrence may be explored, such as sinusoidal or sigmoid activation behavior.
Additionally, the reinitialization is currently bypassed for problems where $N<D$. Testing other redistribution methods, such as uniform quasi-Monte Carlo sequences, may be beneficial for these cases.
7.1.4. Verification:
While QUASAR’s highly stochastic framework makes it a practical choice for non-differentiable and non-parametric problems, the algorithms’s performance would be further validated by rigorously evaluating its performance on constrained or multi-objective problems.